OpenStudy (anonymous):

I have to prove this: (2αsinxcox)^2 +α^2(cos^2 x-sin^2 x)^2= α^2

OpenStudy (freckles):

try using double angle identities to begin with

OpenStudy (freckles):

you should know \[2\sin(x)\cos(x)=\sin(2x) \\ \text{ and } \cos^2(x)-\sin^2(x)=\cos(2x)\]

OpenStudy (campbell_st):

well you a few trig identities 2sin(A)cos(A) = sin(2A) cos^2(A) - sin^2(A) = cos(2A) so you get \[(\alpha \sin(2x))^2 + \alpha^2(\cos(2x))^2 = \alpha^2\] so if you simplify this you get \[\alpha^2\sin^2(2x) + \alpha^2\cos^2(2x) = \alpha^2\] now factor the left hand side and you should see a well known trig identity hope it helps

OpenStudy (anonymous):

\[\sin ^{2}(2x) + \cos ^{2}(2x)= 1 \] just like \[\sin ^{2}x+ \cos ^{2}x=1\] does, right?

OpenStudy (campbell_st):

that's correct